# Efficient Context Selection for Long-Context QA: No Tuning, No Iteration, Just Adaptive-$k$

### v1

@@ Abstract
Retrieval-augmented generation (RAG) and long-context language models (LCLMs) both address context limitations of LLMs in open-domain question answering (QA). However, optimal external context to retrieve remains an open problem: fixing the retrieval size risks either wasting tokens or omitting key evidence. Existing adaptive methods like Self-RAG and Self-Route rely on iterative LLM prompting and perform well on factoid QA, but struggle with aggregation QA, where the optimal context size is both unknown and variable.
We present Adaptive-$k$ retrieval, a simple and effective single-pass method that adaptively selects the number of passages based on the distribution of the similarity scores between the query and the candidate passages. It does not require model fine-tuning, extra LLM inferences or changes to existing retriever–reader pipelines. On both factoid and aggregation QA benchmarks, Adaptive-$k$ matches or outperforms fixed-$k$ baselines while using up to 10× fewer tokens than full-context input, yet still retrieves 70% of relevant passages. It improves accuracy across five LCLMs and two embedding models, highlighting that dynamically adjusting context size leads to more efficient and accurate QA.
@@
**footnotetext: Work done during an internship at Megagon Labs.
@@ 1 Introduction
We evaluate Adaptive-$k$ on both factoid and aggregation-style QA tasks across multiple LCLMs and embedding models. Our experiments span two representative long-context benchmarks: HELMET Yen et al. (2025), which includes factoid QA tasks with up to 128k-token contexts, and HoloBench Maekawa et al. (2025), which focuses on aggregation-style queries. Our results show that on aggregation-QA, Adaptive-$k$ outperforms Self-Route by up to +9 points in answer accuracy on high-information tasks. It consistently maintains $\sim$70context recall and reduces token usage by 2× to 10× compared to full-context baselines. On factoid QA, Adaptive-$k$ matches or exceeds the accuracy of fixed-size retrieval with up to 99% reduction in input tokens, effectively pruning irrelevant content. These findings highlight the importance of query-specific context sizing and establish Adaptive-$k$ as a simple, robust, and efficient alternative to more complex adaptive retrieval strategies.
@@ 3.3 Proposed method
$\bm{C}\leftarrow\texttt{Embedder}(C)$triangleright Precomputed
@@
$\bm{g}\leftarrow\texttt{array}()$triangleright For storing the gap
@@
$k\leftarrow\arg\max(\bm{g})$triangleright Index at the largest gap
@@ 4.1 Dataset
[…]
@@ 4.2 Models
We test our method on small and large embedding models: BAAI’s bge-en-large-v1.5111https://huggingface.co/BAAI/bge-large-en-v1.5 Xiao et al. (2023) and Alibaba NLP’s gte-Qwen2-1.5B-instruct [^2] https://huggingface.co/Alibaba-NLP/gte-Qwen2-1.5B-instruct Li et al. (2023).
@@ 4.3 Compared methods
[…]
fixed-25k (oracle)full-contextSelf-Routeadaptive-k (ours)002020202040404040020k40k60k80k100kScore# Input Tokens
[…]
fixed-25k (oracle)full-contextSelf-Routeadaptive-k (ours)002020202040404040020k40k60k80k100kfixed-25k (oracle)full-contextSelf-Routeadaptive-k (ours)002020202040404040020k40k60k80k100kScore# Input Tokens
[…]
fixed-25k (oracle)full-contextSelf-Routeadaptive-k (ours)002020202040404040020k40k60k80k100kfixed-25k (oracle)full-contextSelf-Routeadaptive-k (ours)002020202040404040020k40k60k80k100kfixed-50k (oracle)full-contextSelf-Routeadaptive-k (ours)002020202040404040020k40k60k80k100kScore# Input Tokens
@@ 5.3 Comparison across LLMs
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 5.4 Embedding bottleneck
Retrieval BGE GTE
HotpotQA Self-Route 90.83 25.83
adaptive-k 70.83 5.50
NQ Self-Route 51.90 20.67
adaptive-k 27.20 2.85
TriviaQA Self-Route 46.52 10.20
adaptive-k 31.21 3.00
HoloBench-5k Self-Route 65.79 65.18
adaptive-k 75.74 82.20
HoloBench-10k Self-Route 45.04 45.87
adaptive-k 68.54 78.99
HoloBench-25k Self-Route 30.42 31.02
adaptive-k 66.16 76.47
HoloBench-50k Self-Route 21.54 21.90
adaptive-k 67.43 72.54
Table 3: A comparison of the context recall scores across tasks between BGE (bge-large-en-v1.5) and GTE (gte-Qwen2-1.5B-instruct).
We observed that the effectiveness of our adaptive method is sensitive to the choice of embedding model. As shown in Table 3, bge-large-en-v1.5 embeddings and gte-Qwen2-1.5B-instruct embeddings have different strengths depending on the task. In factoid QA tasks, BGE embeddings consistently yield higher context recall than GTE, whereas GTE performs better on HoloBench. The underlying cause remains unclear, but we identify a few potential factors: (1) Context chunk length: the factoid QA tasks in HELMET generally have a longer context chunk length (up to $\sim$100 tokens) than HoloBench ($\sim$40 tokens); (2) Chunking scheme Zhong et al. (2025): while the context chunks in HoloBench contain well-formed natural-language sentences, those in the factoid QA tasks often contain mid-sentence breaks; (3) Training scheme: differences in pretraining corpora and formatting may lead to divergent performance across embedding models. Overall, choosing the right embedding model is critical for ensuring RAG effectiveness. For general use, we recommend bge-large-en-v1.5 for Adaptive-$k$ due to its strong and consistent performance across settings.
@@ 6 Conclusion
Our adaptive-$k$ retrieval is a plug-and-play, single-pass alternative to fixed-size retrieval, yet several directions remain. First, because the method is orthogonal to most RAG pipelines, pairing it with techniques such as query-expansion, iterative reranking, or generative feedback loops could further improve accuracy and latency. Second, embedding models excel on different query and corpus traits; a runtime system that selects—or ensembles—embeddings per query may unlock extra gains in recall and robustness.
@@ A.2 Detailed experimental setup
We set temperature and top-p parameters to $0.0$ and $1.0$, respectively, for all our experiments. For Gemini-2.5-Flash, we set its thinking budget to 00. Table 4 lists the models used in our experiments.

### v2

@@ Abstract
Retrieval-augmented generation (RAG) and long-context language models (LCLMs) both address context limitations of LLMs in open-domain question answering (QA). However, optimal external context to retrieve remains an open problem: fixing the retrieval size risks either wasting tokens or omitting key evidence. Existing adaptive methods like Self-RAG and Self-Route rely on iterative LLM prompting and perform well on factoid QA, but struggle with aggregation QA, where the optimal context size is both unknown and variable. We present Adaptive-$k$ retrieval, a simple and effective single-pass method that adaptively selects the number of passages based on the distribution of the similarity scores between the query and the candidate passages. It does not require model fine-tuning, extra LLM inferences or changes to existing retriever–reader pipelines. On both factoid and aggregation QA benchmarks, Adaptive-$k$ matches or outperforms fixed-$k$ baselines while using up to 10× fewer tokens than full-context input, yet still retrieves 70% of relevant passages. It improves accuracy across five LCLMs and two embedding models, highlighting that dynamically adjusting context size leads to more efficient and accurate QA. [^1] The code is available at https://github.com/megagonlabs/adaptive-k-retrieval.
@@ 1 Introduction
We evaluate Adaptive-$k$ on both factoid and aggregation-style QA tasks across multiple LCLMs and embedding models. Our experiments span two representative long-context benchmarks: HELMET Yen et al. (2025), which includes factoid QA tasks with up to 128k-token contexts, and HoloBench Maekawa et al. (2025), which focuses on aggregation-style queries. Our results show that on aggregation-QA, Adaptive-$k$ outperforms Self-Route by up to +9 points in answer accuracy on high-information tasks. It consistently maintains $\sim$70% context recall and reduces token usage by 2× to 10× compared to full-context baselines. On factoid QA, Adaptive-$k$ matches or exceeds the accuracy of fixed-size retrieval with up to 99% reduction in input tokens, effectively pruning irrelevant content. These findings highlight the importance of query-specific context sizing and establish Adaptive-$k$ as a simple, robust, and efficient alternative to more complex adaptive retrieval strategies.
@@ 3.3 Proposed method
$\bm{C}\leftarrow\texttt{Embedder}(C)$ $\triangleright$ Precomputed
@@
$\bm{g}\leftarrow\texttt{array}()$ $\triangleright$ For storing the gap
@@
$k\leftarrow\arg\max(\bm{g})$ $\triangleright$ Index at the largest gap
@@ 4.1 Dataset
For aggregation tasks, we employ HoloBench Maekawa et al. (2025), which provides 90 evaluation samples. HoloBench allows control over both total context size and the amount of information relevant to the query. We fix the total context to 100k tokens and evaluate under varying levels of relevant information, with $\texttt{info\_amount}=$ {5000, 10000, 25000, 50000} tokens.
@@ 4.2 Models
We test our method on small, medium, and large embedding models: Meta’s contriver-msmarco [^2] https://huggingface.co/facebook/contriever-msmarco (Izacard et al., 2021) with 109M params, BAAI’s bge-en-large-v1.5333https://huggingface.co/BAAI/bge-large-en-v1.5 Xiao et al. (2023) with 335M params, and Alibaba NLP’s gte-Qwen2-1.5B-instruct [^4] https://huggingface.co/Alibaba-NLP/gte-Qwen2-1.5B-instruct Li et al. (2023) with 1.78B params.
@@ 4.3 Compared methods
(a) $\texttt{info\_amount}=10000$.
fixed-25k (oracle)full-contextSelf-Routeadaptive-k (ours)020204040020k40k60k80k100kScore# Input Tokens
(b) $\texttt{info\_amount}=25000$.
fixed-25k (oracle)full-contextSelf-Routeadaptive-k (ours)020204040020k40k60k80k100kScore# Input Tokens
(c) $\texttt{info\_amount}=50000$.
fixed-50k (oracle)full-contextSelf-Routeadaptive-k (ours)020204040020k40k60k80k100kScore# Input Tokens
@@ 5.3 Comparison across LLMs
fixed-25k (oracle)full-contextSelf-Routeadaptive-k (ours)0202040406060020k40k60k80k100kScore# Input Tokens
@@
fixed-25k (oracle)full-contextSelf-Routeadaptive-k (ours)0202040406060020k40k60k80k100kScore# Input Tokens
@@
fixed-50k (oracle)full-contextSelf-Routeadaptive-k (ours)0202040406060020k40k60k80k100kScore# Input Tokens
@@
fixed-5k (oracle)full-contextSelf-Routeadaptive-k (ours)0202040406060020k40k60k80k100kScore# Input Tokens
@@
fixed-50k (oracle)full-contextSelf-Routeadaptive-k (ours)0202040406060020k40k60k80k100kScore# Input Tokens
@@ 5.4 Embedding bottleneck
Method Contriever BGE GTE
HotpotQA Self-Route 77.50 90.83 25.83
adaptive-k 49.17 70.83 5.50
NQ Self-Route 51.19 51.90 20.67
adaptive-k 26.28 27.20 2.85
TriviaQA Self-Route 41.49 46.52 10.20
adaptive-k 25.02 31.21 3.00
HoloBench-5k Self-Route 60.57 65.79 65.18
adaptive-k 72.11 75.74 82.20
HoloBench-10k Self-Route 41.83 45.04 45.87
adaptive-k 70.85 68.54 78.99
HoloBench-25k Self-Route 29.67 30.42 31.02
adaptive-k 68.67 66.16 76.47
HoloBench-50k Self-Route 21.27 21.54 21.90
adaptive-k 67.16 67.43 72.54
Table 3: A comparison of the context recall scores across tasks between Contriever (contriever-msmarco), BGE (bge-large-en-v1.5), and GTE (gte-Qwen2-1.5B-instruct).
We observed that the effectiveness of our adaptive method is sensitive to the choice of embedding model. As shown in Table 3, the embeddings by bge-large-en-v1.5, gte-Qwen2-1.5B-instruct, and contriever-msmarco have different strengths depending on the task. In factoid QA tasks, BGE embeddings consistently yield higher context recall than GTE, whereas GTE performs better on HoloBench. The underlying cause remains unclear, but we identify a few potential factors: (1) Context chunk length: the factoid QA tasks in HELMET generally have a longer context chunk length (up to $\sim$100 tokens) than HoloBench ($\sim$40 tokens); (2) Chunking scheme Zhong et al. (2025): while the context chunks in HoloBench contain well-formed natural-language sentences, those in the factoid QA tasks often contain mid-sentence breaks; (3) Training scheme: differences in pretraining corpora and formatting may lead to divergent performance across embedding models. Overall, choosing the right embedding model is critical for ensuring RAG effectiveness. For general use, we recommend bge-large-en-v1.5 for Adaptive-$k$ due to its strong and consistent performance across settings.
@@ 5.5 Limitation of fixed retrieval
In addition, Tables 4 and 5 demonstrate that the Adaptive-$k$ method effectively estimates the threshold of relevant contexts across different relevant context sizes. These tables report the average absolute difference between the estimated threshold ($k$) and the true $k$-value, referred to as diff-$k$ henceforth, across various datasets and embeddings. The true $k$-value is defined as the index of the last relevant context chunk in the list of context chunk embeddings sorted by similarity; in other words, diff-$k$ is the smallest value of $k$ that achieves 100% recall. Table 4 presents the diff-$k$ values for the HoloBench task with 50k tokens of relevant context. The diff-$k$ values for Adaptive-$k$ are substantially lower than those of the Self-Route baseline and closely match those of the fixed 50k-token retrieval, which is treated as an oracle. Similarly, the diff-$k$ values for Adaptive-$k$ in the HotpotQA task (Table 5) consistently approximate the oracle retrieval results, indicating that the method reliably determines the optimal retrieval amount.
Contriever BGE GTE
fixed-1k 2351.96 2289.33 2265.12
fixed-5k 2236.31 2186.08 2154.81
fixed-10k 2095.38 2056.89 2017.09
fixed-25k 1680.93 1660.64 1599.99
fixed-50k (oracle) 992.47 985.39 898.34
Adaptive-$k$ 1133.28 1248.99 1001.00
Self-Route 2236.31 2186.09 2154.81
Table 4: Diff-$k$ values in the HoloBench task with an information amount of 50k for the three embedding models.
Contriever BGE GTE
fixed-1k (oracle) 162.52 88.86 631.90
fixed-5k 154.68 97.89 598.70
fixed-10k 156.29 116.35 558.69
fixed-25k 191.14 185.70 452.32
fixed-50k 297.41 320.50 307.12
Adaptive-$k$ 161.14 89.40 632.28
Self-Route 154.68 97.89 598.70
Table 5: Diff-$k$ values in HotpotQA for the three embedding models. Given the low retrieval performance with the GTE embeddings in the factoid-type QA tasks in Table 3, the results of GTE embeddings in this table are not informative.
@@ 6 Conclusion
Our adaptive-$k$ retrieval is a plug-and-play, single-pass alternative to fixed-size retrieval, yet several directions remain. First, because the method is orthogonal to most RAG pipelines, pairing it with techniques such as query-expansion, iterative reranking, or generative feedback loops could further improve accuracy and latency. Second, embedding models excel on different query and corpus traits; a runtime system that selects or ensembles embeddings per query may unlock extra gains in recall and robustness.
@@
Acknowledgments
We thank Hayate Iso and Pouya Pezeshkpour for their constructive feedback on this study.
@@ A.2 Detailed experimental setup
We set temperature and top-p parameters to $0.0$ and $1.0$, respectively, for all our experiments. For Gemini-2.5-Flash, we set its thinking budget to 0. Table 6 lists the models used in our experiments.